OpenStudy (anonymous):

Select the equation in which the graph of the line has a negative slope, and the y-intercept equals -4. 5x + 3y = 4 5x - 3y =-4 4x + 16y = -20 6x + 2y = -8

pooja195 (pooja195):

ok for each equation do the X and Y intercept do you know how?

pooja195 (pooja195):

\[5x + 3(0) = 4\] \[5(0) + 3y = 4\]

OpenStudy (anonymous):

no, i really don't understand any of this, i am super behind in algebra and thought i remembered this stuff but i was sorely mistaken : ((

pooja195 (pooja195):

ok solve the equations i gave :)

OpenStudy (anonymous):

I see hows its done.

OpenStudy (anonymous):

okay, the first equation is ... 5x+3(0)=4 ... so first i multiply 3 * 0, which equals 0 of course. then the equation is 5x + 0 = 4, so we simplify it to 5x = 4. The second equation is ... 5(0)+3y=4 ... so that equation turns out to be 3y = 4 after we do the steps above again. is that correct ?

pooja195 (pooja195):

im not sure...anymore ;-; im gonna let someone else take this one this is confusing,,,

pooja195 (pooja195):

@sleepyjess @TheSmartOne

OpenStudy (anonymous):

thats fine, i understand ! hopefully someone else can help

OpenStudy (anonymous):

@pooja195 I think you are doing it right with the two equation things.

OpenStudy (anonymous):

Idk either. But the equation part is correct i think. :)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

ill try

pooja195 (pooja195):

if thats the case then \[3y = 4 \] divide by 3 so 4/3 and then same with the other equation \[5x=4\] divide by 5 so 5/4

OpenStudy (sleepyjess):

Right now, let's just use y = mx + b, then convert it to standard form. We know that the slope is negative, so -m, and that the y intercept is -4. So that gives us y = -mx - 4

OpenStudy (sleepyjess):

Maybe... this is confusing. I know the answer, but I don't know how to explain. Or at least I think I know the answer

pooja195 (pooja195):

@amistre64 do you know?

OpenStudy (amistre64):

when x=0, y=-4 is a solution that we can narrow the options with

OpenStudy (amistre64):

theres only one option that satisfies this criteria

OpenStudy (anonymous):

... 5x - 3y = -4 ?? :oo

OpenStudy (radar):

Naw, I don't think that one is it.....because rearranging it you get 3y = 5x + 4 no -4, no negative slope, can't be.

OpenStudy (anonymous):

Hmmm.. This is hard !!! Grr

OpenStudy (radar):

The post by @sleepyjess tells you how to solve it, amistre64 shows a practical way to spot the correct solution. There is only one when placed in the slope intercept form has an intercept of -4.

OpenStudy (sleepyjess):

Oh, so I was on the right track...

OpenStudy (amistre64):

there were tracks?

OpenStudy (radar):

yes, wise-tracks lol

OpenStudy (sleepyjess):

lol

OpenStudy (radar):

OK harley_may looking at the first two can you see we can eliminate them because the intercept is not a negative ?

OpenStudy (anonymous):

Yes, I think I get that

OpenStudy (radar):

those would be positive as well as improper value. Now look at the third one...what do you think. 4x + 16y = -20

OpenStudy (radar):

First for kicks divide thru by 4 and look at it again.

OpenStudy (radar):

Do you see: x + 4y = -5 intercept is negative but alas it is not a -4.

OpenStudy (anonymous):

ohhh, i get that ! okay

OpenStudy (radar):

finally look at the last one, 6x + 2y = -8 lets follow sleepyjess suggestion 2y = -6x -8 y=-3x - 4 in slope intercept form. what you think?

OpenStudy (anonymous):

i see a negative four in there !! is that the answer ?

OpenStudy (radar):

Yes, it has a negative slope (-6) and an intercept of -4 A perfect candidate .

OpenStudy (anonymous):

thank you so so so much !!!

OpenStudy (radar):

You're welcome, and wishing you good luck with your studies.